0 0
Add a comment Improve this question Transcribed image text
Answer #1

A

Let the no. of people sampled be x.
p = 16%

We need the expected no. of people to be atleast 25.
Thus, x*0.16 = 25
x = 156.25
Hence, we need to sample 157 people.

B

Let the no. of people sampled be x.
p = 12%

We need the expected no. of people to be atleast 25.
Thus, x*0.12 = 25
x = 208.33
Hence, we need to sample 209 people.

C

Let the no. of people sampled be x.
p = 2%

We need the expected no. of people to be atleast 25.
Thus, x*0.02 = 25
x = 1250
Hence, we need to sample 1250 people.

D

Std Dev = [157*0.16*(1-0.16)]0.5 = 4.5936

E

Std Dev = [209*0.12*(1-0.12)]0.5 = 4.6979
Answer #2

A

Let the no. of people sampled be x.
p = 16%

We need the expected no. of people to be atleast 25.
Thus, x*0.16 = 25
x = 156.25
Hence, we need to sample 157 people.

B

Let the no. of people sampled be x.
p = 12%

We need the expected no. of people to be atleast 25.
Thus, x*0.12 = 25
x = 208.33
Hence, we need to sample 209 people.

C

Let the no. of people sampled be x.
p = 2%

We need the expected no. of people to be atleast 25.
Thus, x*0.02 = 25
x = 1250
Hence, we need to sample 1250 people.

D

Std Dev = [157*0.16*(1-0.16)]0.5 = 4.5936

E

Std Dev = [209*0.12*(1-0.12)]0.5 = 4.6979